How much interest will you earn on a $792 deposit at 6% compound semiannually if you deposit for 7 years?
Answer #1

This question requires application of basic time value of money function, according to which FV = PV * (1 + r) where FV is F

r = 6%/2 = 3% (semi-annually), n = 7 * 2 = 14 semi-annual periods

FV = $792 * (1 + 0.03)14

FV = $792 * 1.51

FV = $1,197.97

Interest = $1,197.97 - $792 = $405.97